Suppose that (π‘₯∝)βˆπœ–π½ ⟢ π‘₯ in X and (π‘¦βˆ)βˆπœ–π½ ⟢ 𝑦 in Y. Show that

(π‘₯𝛼 Γ— 𝑦𝛼) β†’ π‘₯ Γ— 𝑦 𝑖𝑛 𝑋 Γ— π‘Œ.


1
Expert's answer
2020-09-24T11:57:21-0400

(π‘₯𝛼 Γ— 𝑦𝛼) β†’ π‘₯ Γ— 𝑦 𝑖𝑛 𝑋 Γ— π‘Œ


proof considering the limits

lim "\alpha\to1" "(x^\alpha)^\alpha = X" and lim "\alpha""\to1" "(y^\alpha)^\alpha = Y"


lim "\alpha""\to1" "(x^\alpha)^\alpha = \alpha x ," lim "\alpha""\to1" "(y^\alpha)^\alpha = \alpha Y"


since "\alpha\to1" then (π‘₯𝛼 Γ— 𝑦𝛼) β†’ π‘₯ Γ— 𝑦 𝑖𝑛 𝑋 Γ— π‘Œ